OpenStudy (anonymous):

given points R (-2,3) and S (5,1) find the slope of every line parallel to line RS A.slope of line RS B.slope of perpendicular lines

OpenStudy (anonymous):

the slop of RS m= (1-3)/(5-(-2))=-2/7

OpenStudy (anonymous):

the slope of lines parallel to it is the same slope m = -2/7

OpenStudy (anonymous):

the slope of the perpendicular lines is the negative reciprocal of m, (call it m2) m2=-1/m=7/2

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, then you should be a fan :P